# Driving with DINO: Vision Foundation Features as a Unified Bridge for Sim-to-Real Generation in Autonomous Driving

### v1

@@ Abstract.
diffusion models, video Sim2Real
copyright: none
Figure 1. Driving with DINO (DwD) achieves photorealistic simulation-to-real video translation with superior structural consistency. As shown, our method generalizes remarkably well across diverse scenarios, ranging from town and urban environments to challenging weather conditions such as rain and frost.
1. Introduction
Closed-loop validation is critical for autonomous driving safety. However, applying it in the real world is risky and inefficient due to the massive data requirements and the scarcity of rare corner cases. Simulation offers a scalable alternative, yet creating photorealistic assets remains costly. Consequently, simulators like CARLA (dosovitskiy2017carla) often compromise on visual fidelity, resulting in a significant domain gap between simulated and real-world data. This discrepancy hinders the generalization of simulation-trained models to physical environments, necessitating robust Sim-to-Real transfer techniques to effectively utilize simulation assets.
Figure 2. Consistency-Realism-Dilemma
To address this challenge, Diffusion Models have emerged as a promising paradigm, offering superior stability and high-fidelity generation capabilities. While initial attempts (meng2021sdedit; parmar2024onestep) adapted these image diffusion models for Sim-to-Real in a frame-wise manner; however, treating continuous driving video as independent images ignores temporal correlations, inevitably leading to severe flickering artifacts and texture incoherence. To remedy this, leveraging video diffusion models augmented with structural control mechanisms (e.g., ControlNet (zhang2023adding)) emerges as a superior paradigm (zhang2023controlvideo; wang2023videocomposer; guo2024sparsectrl; he2024cameractrl; ma2024trailblazer; xi2025omnivdiff; jiang2025vace). Video diffusion inherently model temporal dynamics to ensure coherence, while the integration of ControlNet adapters strictly anchors the generative process to the simulation’s geometric layout. This integration effectively reconciles the tension between temporal and spatial consistency, making it a highly promising direction for robust Sim-to-Real transfer.
In this context, intermediate representations such as edge maps, blurred images, semantic masks, and depth maps serve as the common bridge between simulation and reality, and have been widely integrated into ControlNet-based frameworks (alhaija2025cosmos; ali2025world; gao2023magicdrive; gao2025magicdrive). Additionally, High-Definition Maps (HDMaps) is a widely adopted structural representation in the autonomous driving domain. However, achieving both high photorealism and spatial consistency simultaneously remains a significant challenge, which we call as Consistency-Realism-Dilemma. As depicted in Fig. 2, current widely adopted control signals could not balance them well:
@@ 1. Introduction
To leverage the advantages of both feature types, recent multi-control frameworks attempt to mitigate these trade-offs, yet they face significant challenges. Jointly training multiple ControlNets incurs prohibitive GPU memory costs and suffers from gradient conflicts, where heterogeneous modalities compete and fail to converge (chen2023control; hu2023videocontrolnet). Conversely, combining separately trained adapters during inference is equally problematic. Aggregating distinct control signals often leads to feature interference, necessitating laborious weight tuning and frequently resulting in incoherent fusion or visual artifacts due to conflicting guidance (Zhao2023).
Departing from complex multi-branch designs, we propose a novel conditioning strategy leveraging Vision Foundation Models (VFMs), specifically DINOv3 (simeoni2025dinov3). Unlike explicit multi-control approaches, DINO features inherently encode a rich spectrum of scene properties ranging from low-level to high-level cues within a unified latent space, a capability validated by its robust performance across a wide spectrum of task granularities (kirillov2023segment; wang2025vggt; wang2025pi; yang2024depth; xu2025pixel; zheng2025diffusion). By utilizing DINO as a holistic condition for Sim-to-Real transfer, we achieve a natural fusion of structural and semantic information. Crucially, this approach eliminates the feature competition and gradient conflicts typical of multi-branch architectures, allowing for efficient training and inference without additional computational overhead or complex manual tuning.
Nevertheless, our analysis reveals that naively incorporating DINO features into ControlNet results in textural information leakage. Since DINO representations are rich enough to serve as powerful encoders for auto-encoding tasks—a property explored in prior work (zheng2025diffusion; jia2025dino; shi2025latent)—they inadvertently carry over fine-grained pixel information. In our Sim-to-Real context, the model essentially memorizes and reproduces the synthetic textures of the simulation instead of generating realistic details, thereby defeating the purpose of domain adaptation. Furthermore, the spatial resolution of DINO features is typically down-sampled by a factor of 16. This significant compression results in the loss of critical structural details, particularly along object boundaries, which inevitably degrades control consistency. Conversely, in the temporal domain, DINO features exhibit significant redundancy due to their independent frame-wise processing nature. While existing methods often resort to naive keyframe sampling to mitigate this (burgert2025go; russell2025gaia), our analysis suggests that such aggressive subsampling disrupts motion continuity, inevitably leading to temporal aliasing and incoherence in the generated video.
@@
Figure 3. The Training and Inference pipeline of DwD.
2. Related Works
2.1. Generative Adversarial Networks
@@
Early Sim-to-Real approaches utilized GANs (e.g., CycleGAN (zhu2017unpaired), CUT (park2020contrastive)) to align simulated images with real-world distributions. However, these adversarial translations frequently introduce geometric distortions and semantic inconsistencies. While utilizing simulator-derived G-buffers (e.g., depth, normals) can constrain the generation process to preserve geometry (richter2022enhancing), such methods often retain adversarial artifacts and lack fine-grained texture fidelity.
2.2. 2D Image and Video Diffusion
@@
Diffusion models have mitigated GAN-based instability through superior fine-grained fidelity. SDEdit (meng2021sdedit) facilitates transfer via SDE inversion (song2020denoising), injecting photorealistic priors while maintaining the simulator’s semantic structure. Extensions to video (e.g., Pix2Video (ceylan2023pix2video), TokenFlow (geyer2023tokenflow)) enforce temporal coherence within zero-shot frameworks by propagating features or injecting optical flow priors (FLATTEN (cong2023flatten)).
@@
However, methods relying on inversion or diffusion-based relighting (zhang2025scaling; liu2025tc) struggle to effectively disentangle semantic structure from synthetic style, resulting in severe “texture baking.” To address this, ControlNet (zhang2023adding) introduced explicit spatial conditioning to decouple geometry from appearance. With the rise of Video Diffusion Transformers (DiTs) (peebles2023scalable), controllable generation has shifted towards inherently modeling motion dynamics, offering temporal consistency superior to zero-shot adaptations.
2.3. Controllable Video Diffusion
@@
In autonomous driving simulation, methods like MagicDrive (gao2023magicdrive; gao2025magicdrive) and DriveDreamer (wang2024drivedreamer; zhao2025drivedreamer) utilize sparse high-level priors (e.g., HDMaps, 3D boxes). While effective for layout, the spatial sparsity of these signals leads to hallucinations in unconstrained regions. Conversely, volumetric approaches like WoVoGen (lu2024wovogen) utilize 3D occupancy for consistency but remain too coarse-grained for fine geometric details.
@@
Recent efforts to unify these paradigms, such as Cosmos Transfer (alhaija2025cosmos; ali2025world), incorporate both high-level semantic and low-level geometric cues. However, fusing such heterogeneous signals introduces significant architectural complexity and instability, often hitting a Consistency-Realism Dilemma. Existing multi-control fusion strategies are either limited to similar modalities (gao2025longvie) or suffer from prohibitive training costs (xi2025ctrlvdiff). To address these limitations, we propose utilizing DINOv3 as a unified representation that encapsulates both semantics and structures, allowing it to be seamlessly integrated into a standard single-control architecture. While DINO has been applied to object-level control (chen2024anydoor; jiang2025dive), we extend this to scene-level Sim-to-Real transfer, utilizing a unified representation to reconcile photorealism with strict structural consistency and eliminate texture baking.
@@
Figure 4. Lower-dimensional PCA components encode coarse semantic layouts, whereas increasing the dimensions leads to the refinement of high-frequency details. Similarity maps are shown for two anchor points (top: $\times$ on building, bottom: $\times$ on one car).
3. Method
3.1. Overview
@@
3.2. Controllable Video Diffusion Model
The DwD framework leverages the architecture of the pretrained Cosmos-Predict2.5 (ali2025world), which integrates the WAN2.1 3D VAE (wan2025wan) with fine-tuned Diffusion Transformer (DiT) blocks. To enable structurally controllable generation, we augment the video diffusion framework with a dedicated control branch, as illustrated in Fig. 3(a). This branch comprises conditioning DiT blocks initialized with weights from the denoising DiT backbone. Following the injection strategy proposed in VACE (jiang2025vace), control features are modulated into the base model at uniform intervals of $N$ blocks. The control branch encodes processed DINOv3 latents as conditioning inputs. During the training phase, the model is optimized to reconstruct ground-truth realistic videos. At inference( Fig. 3(b), the model processes synthetic simulation inputs to generate photorealistic counterparts, effectively bridging the domain gap between simulation and reality.
3.3. Robust Training with VFM-Prism
@@
Spatial Resolution Enhancer. Unlike previous VFMs (radford2021learning; oquab2023dinov2; tschannen2025siglip) that exhibit bias against high-resolution inputs, DINOv3 demonstrates robustness to variable resolutions due to its specialized positional encoding strategy (RoPE-box jittering). To compensate for the high spatial compression rate of DINOv3, we upsample the input video $\boldsymbol{V}\in\mathbb{R}^{T\times H\times W\times 3}$ by a scale factor $S$ prior to encoding, yielding a higher-resolution feature map $\boldsymbol{Z}_{c}\in\mathbb{R}^{T\times\frac{H\times S}{16}\times\frac{W\times S}{16}\times C}$. While post-hoc feature map upsampling methods like Featup (fu2024featup) and Anyup (wimmer2025anyup) offer computationally cheaper alternatives, our empirical findings suggest they compromise feature map quality, negatively impacting the consistency between generated videos and input simulations. Thus, input-space upscaling ensures superior structural fidelity. However, this results in a latent tensor size misaligned with the input requirements of the Condition DiT blocks, which should spatially be $\frac{H}{16}\times\frac{W}{16}$ rather than $\frac{H\times S}{16}\times\frac{W\times S}{16}$. Therefore, we introduce a Spatial Alignment Module to bridge this gap, which is composed of strided convolutions with residual blocks, as detailed in the Appendix.
@@
Causal Temporal Aggregator. The absence of temporal compression in DINOv3, results in a temporal dimension misalignment with the input requirements of the Condition DiT blocks. While prior methods rely on heuristic downsampling strategies such as keyframe sampling (burgert2025go; hwang2025cross) or bilinear interpolation (russell2025gaia), these approaches often fail to capture fine-grained spatial details and temporal dynamics. To bridge this gap without incurring the computational overhead of a full CausalVAE, we insert extra causal convolution blocks into aforementioned Spatial Alignment Module (see Appendix).
3.4. Video Sim2Real Translation
@@
4. Experiments
4.1. Experiment Setup
We train DwD on the nuPlan dataset (caesar2021nuplan) and evaluate performance using videos rendered from the CARLA simulator (see Appendix for more details). To ensure a comprehensive assessment, we compare our method against representative baselines across different categories. For training-free image diffusion models like TC-Light (liu2025tc), we adhere to the official protocols using appropriate relighting prompts. For FRESCO (yang2024fresco), we additionally incorporate nuPlan images as reference style inputs. In the video generation domain, we compare against Cosmos-Transfer2.5 (ali2025world) with various input modalities. Crucially, to ensure a fair comparison, we fine-tune the ControlNet modules of these video generation baselines on the same training set used for our method.
Figure 5. Qualitative comparison with state-of-the-art methods. The red boxes highlight inconsistencies with respect to the CG input, while the green boxes point out low-fidelity textures.
4.2. Evaluation Metrics.
@@ 4.1. Experiment Setup
  • Visual Fidelity: We employ sFID and sKID to measure distribution discrepancy, adopting a semantics-aware sampling strategy (richter2022enhancing) to mitigate layout mismatches. By matching VGG features between CG and real patches, we construct a benchmark of 400,000 pairs for robust evaluation.
  • Perceptual Realism: We quantify authenticity via CLIP-Real (zeng2025neuralremaster), formulated as $\text{CLIP-Real}=(x^{\top}t_{p})/(x^{\top}t_{n})$. Here, $x,t_{p},t_{n}$ represent embeddings of the output frame, positive (e.g., “Photo”), and negative (e.g., “Game”) prompts, respectively. Higher scores indicate better alignment with real photography.
  • Temporal Consistency: We assess temporal coherence using Motion-S (huang2024vbench) for motion plausibility and WarpSSIM. The latter computes SSIM between a frame and its neighbors warped via optical flow derived from source CG videos.
  • Sim2Real Consistency: We assess semantic preservation using the mean Intersection over Union (mIoU). Specifically, we employ the state-of-the-art perception model (wang2023internimage) (pretrained on Cityscapes) to perform semantic segmentation on the translated frames. We then calculate the mIoU by comparing the predicted segmentation masks against the ground truth semantic labels derived from the source CG engine.
4.3. Quantitative Evaluation.
@@
Table 1. Quantitative evaluations. Top results are colored in Gold, Silver, and Bronze(also applicable to Tabs. 2 and 3.). $\dagger$ denotes modules fine-tuned on the same dataset as DwD. C-T2.5 stands for Cosmos-Transfer 2.5.
@@
[…]
C-T2.5 Edge+Depth $\cellcolorabc93.97$ 112.15 23.71 36.69
[…]
4.4. Qualitative Evaluation.

### v2

@@ Abstract
1 Introduction
Closed-loop validation is critical for autonomous driving safety. However, applying it in the real world is risky and inefficient due to the massive data requirements and the scarcity of rare corner cases. Simulation offers a scalable alternative, yet creating photorealistic assets remains costly. Consequently, simulators like CARLA [10] often compromise on visual fidelity, resulting in a significant domain gap between simulated and real-world data. This discrepancy hinders the generalization of simulation-trained models to physical environments, necessitating robust Sim-to-Real transfer techniques to effectively utilize simulation assets.
Figure 1: Consistency-Realism-Dilemma
To address this challenge, Diffusion Models have emerged as a promising paradigm, offering superior stability and high-fidelity generation capabilities. While initial attempts  [28, 31] adapted these image diffusion models for Sim-to-Real in a frame-wise manner; however, treating continuous driving video as independent images ignores temporal correlations, inevitably leading to severe flickering artifacts and texture incoherence. To remedy this, leveraging video diffusion models augmented with structural control mechanisms (e.g., ControlNet [53]) emerges as a superior paradigm [55, 43, 16, 17, 27, 48, 23]. Video diffusion inherently model temporal dynamics to ensure coherence, while the integration of ControlNet adapters strictly anchors the generative process to the simulation’s geometric layout. This integration effectively reconciles the tension between temporal and spatial consistency, making it a highly promising direction for robust Sim-to-Real transfer.
In this context, intermediate representations such as edge maps, blurred images, semantic masks, and depth maps serve as the common bridge between simulation and reality, and have been widely integrated into ControlNet-based frameworks [1, 2, 14, 13]. Additionally, High-Definition Maps (HDMaps) is a widely adopted structural representation in the autonomous driving domain. However, achieving both high photorealism and spatial consistency simultaneously remains a significant challenge, which we call as Consistency-Realism-Dilemma. As depicted in Fig. 1, current widely adopted control signals could not balance them well:
@@ 1 Introduction
To leverage the advantages of both feature types, recent multi-control frameworks attempt to mitigate these trade-offs, yet they face significant challenges. Jointly training multiple ControlNets incurs prohibitive GPU memory costs and suffers from gradient conflicts, where heterogeneous modalities compete and fail to converge  [7, 18]. Conversely, combining separately trained adapters during inference is equally problematic. Aggregating distinct control signals often leads to feature interference, necessitating laborious weight tuning and frequently resulting in incoherent fusion or visual artifacts due to conflicting guidance  [57].
Departing from complex multi-branch designs, we propose a novel conditioning strategy leveraging Vision Foundation Models (VFMs), specifically DINOv3  [37]. Unlike explicit multi-control approaches, DINO features inherently encode a rich spectrum of scene properties ranging from low-level to high-level cues within a unified latent space, a capability validated by its robust performance across a wide spectrum of task granularities [24, 41, 45, 50, 49, 58]. By utilizing DINO as a holistic condition for Sim-to-Real transfer, we achieve a natural fusion of structural and semantic information. Crucially, this approach eliminates the feature competition and gradient conflicts typical of multi-branch architectures, allowing for efficient training and inference without additional computational overhead or complex manual tuning.
Nevertheless, our analysis reveals that naively incorporating DINO features into ControlNet results in textural information leakage. Since DINO representations are rich enough to serve as powerful encoders for auto-encoding tasks—a property explored in prior work [58, 21, 36]—they inadvertently carry over fine-grained pixel information. In our Sim-to-Real context, the model essentially memorizes and reproduces the synthetic textures of the simulation instead of generating realistic details, thereby defeating the purpose of domain adaptation. Furthermore, the spatial resolution of DINO features is typically down-sampled by a factor of 16. This significant compression results in the loss of critical structural details, particularly along object boundaries, which inevitably degrades control consistency. Conversely, in the temporal domain, DINO features exhibit significant redundancy due to their independent frame-wise processing nature. While existing methods often resort to naive keyframe sampling to mitigate this [4, 35], our analysis suggests that such aggressive subsampling disrupts motion continuity, inevitably leading to temporal aliasing and incoherence in the generated video.
@@
Figure 2: The framework of DwD. (a) Training: The model is trained on real-world driving videos using a controllable diffusion architecture. The core module, VFM-Prism, processes DINOv3 features through Spatial Resolution Enhancement, Minor Components Pruning, and Causal Temporal Aggregation. These refined features are injected via a Control Branch to guide the reconstruction of the original video. (b) Inference: The model performs Sim-to-Real translation using synthetic inputs. By leveraging the domain-invariant structural features extracted by VFM-Prism (specifically via PCA-based pruning to mitigate texture leakage), DwD generates high-fidelity photorealistic videos that strictly preserve the simulation’s geometric layout.
2 Related Works
2.1 Generative Adversarial Networks
@@
Early Sim-to-Real approaches utilized GANs (e.g., CycleGAN [59], CUT [30]) to align simulated images with real-world distributions. However, these adversarial translations frequently introduce geometric distortions and semantic inconsistencies. While utilizing simulator-derived G-buffers (e.g., depth, normals) can constrain the generation process to preserve geometry [34], such methods often retain adversarial artifacts and lack fine-grained texture fidelity.
2.2 2D Image and Video Diffusion
@@
Diffusion models have mitigated GAN-based instability through superior fine-grained fidelity. SDEdit [28] facilitates transfer via SDE inversion [38], injecting photorealistic priors while maintaining the simulator’s semantic structure. Extensions to video (e.g., Pix2Video [6], TokenFlow [15]) enforce temporal coherence within zero-shot frameworks by propagating features or injecting optical flow priors (FLATTEN [9]).
@@
However, methods relying on inversion or diffusion-based relighting [54, 25] struggle to effectively disentangle semantic structure from synthetic style, resulting in severe “texture baking.” To address this, ControlNet [53] introduced explicit spatial conditioning to decouple geometry from appearance. With the rise of Video Diffusion Transformers (DiTs) [32], controllable generation has shifted towards inherently modeling motion dynamics, offering temporal consistency superior to zero-shot adaptations.
2.3 Controllable Video Diffusion
@@
In autonomous driving simulation, methods like MagicDrive [14, 13] and DriveDreamer [44, 56] utilize sparse high-level priors (e.g., HDMaps, 3D boxes). While effective for layout, the spatial sparsity of these signals leads to hallucinations in unconstrained regions. Conversely, volumetric approaches like WoVoGen [26] utilize 3D occupancy for consistency but remain too coarse-grained for fine geometric details.
@@
Recent efforts to unify these paradigms, such as Cosmos Transfer [1, 2], incorporate both high-level semantic and low-level geometric cues. However, fusing such heterogeneous signals introduces significant architectural complexity and instability, often hitting a Consistency-Realism Dilemma. Existing multi-control fusion strategies are either limited to similar modalities [12] or suffer from prohibitive training costs [47]. To address these limitations, we propose utilizing DINOv3 as a unified representation that encapsulates both semantics and structures, allowing it to be seamlessly integrated into a standard single-control architecture. While DINO has been applied to object-level control [8, 22], we extend this to scene-level Sim-to-Real transfer, utilizing a unified representation to reconcile photorealism with strict structural consistency and eliminate texture baking.
@@
Figure 3: Lower-dimensional PCA components encode coarse semantic layouts, whereas increasing the dimensions leads to the refinement of high-frequency details. Similarity maps are shown for two anchor points (top: $\times$ on building, bottom: $\times$ on one car).
3 Method
3.1 Overview
@@
3.2 Controllable Video Diffusion Model
The DwD framework leverages the architecture of the pretrained Cosmos-Predict2.5 [2], which integrates the WAN2.1 3D VAE [40] with fine-tuned Diffusion Transformer (DiT) blocks. To enable structurally controllable generation, we augment the video diffusion framework with a dedicated control branch, as illustrated in Fig. 2(a). This branch comprises conditioning DiT blocks initialized with weights from the denoising DiT backbone. Following the injection strategy proposed in VACE [23], control features are modulated into the base model at uniform intervals of $N$ blocks. The control branch encodes processed DINOv3 latents as conditioning inputs. During the training phase, the model is optimized to reconstruct ground-truth realistic videos. At inference( Fig. 2(b), the model processes synthetic simulation inputs to generate photorealistic counterparts, effectively bridging the domain gap between simulation and reality.
3.3 Robust Training with VFM-Prism
@@
Spatial Resolution Enhancer. Unlike previous VFMs [33, 29, 39] that exhibit bias against high-resolution inputs, DINOv3 demonstrates robustness to variable resolutions due to its specialized positional encoding strategy (RoPE-box jittering). To compensate for the high spatial compression rate of DINOv3, we upsample the input video $\boldsymbol{V}\in\mathbb{R}^{T\times H\times W\times 3}$ by a scale factor $S$ prior to encoding, yielding a higher-resolution feature map $\boldsymbol{Z}_{c}\in\mathbb{R}^{T\times\frac{H\times S}{16}\times\frac{W\times S}{16}\times C}$. While post-hoc feature map upsampling methods like Featup [11] and Anyup [46] offer computationally cheaper alternatives, our empirical findings suggest they compromise feature map quality, negatively impacting the consistency between generated videos and input simulations. Thus, input-space upscaling ensures superior structural fidelity. However, this results in a latent tensor size misaligned with the input requirements of the Condition DiT blocks, which should spatially be $\frac{H}{16}\times\frac{W}{16}$ rather than $\frac{H\times S}{16}\times\frac{W\times S}{16}$. Therefore, we introduce a Spatial Alignment Module to bridge this gap, which is composed of strided convolutions with residual blocks, as detailed in the Appendix.
@@
Causal Temporal Aggregator. The absence of temporal compression in DINOv3, results in a temporal dimension misalignment with the input requirements of the Condition DiT blocks. While prior methods rely on heuristic downsampling strategies such as keyframe sampling [4, 20] or bilinear interpolation [35], these approaches often fail to capture fine-grained spatial details and temporal dynamics. To bridge this gap without incurring the computational overhead of a full CausalVAE, we insert extra causal convolution blocks into aforementioned Spatial Alignment Module (see Appendix).
3.4 Video Sim2Real Translation
@@
4 Experiments
4.1 Experiment Setup
We train DwD on the nuPlan dataset [5] and evaluate performance using videos rendered from the CARLA simulator (see Appendix for more details). To ensure a comprehensive assessment, we compare our method against representative baselines across different categories. For training-free image diffusion models like TC-Light [25], we adhere to the official protocols using appropriate relighting prompts. For FRESCO [51], we additionally incorporate nuPlan images as reference style inputs. In the video generation domain, we compare against Cosmos-Transfer2.5 [2] with various input modalities. Crucially, to ensure a fair comparison, we fine-tune the ControlNet modules of these video generation baselines on the same training set used for our method.
Figure 4: Qualitative comparison with state-of-the-art methods. The red boxes highlight inconsistencies with respect to the CG input, while the green boxes point out low-fidelity textures.
4.2 Evaluation Metrics.
@@ 4.1 Experiment Setup
  • Visual Fidelity: We employ sFID and sKID to measure distribution discrepancy, adopting a semantics-aware sampling strategy [34] to mitigate layout mismatches. By matching VGG features between CG and real patches, we construct a benchmark of 400,000 pairs for robust evaluation.
  • Perceptual Realism: We quantify authenticity via CLIP-Real [52], formulated as $\text{CLIP-Real}=(x^{\top}t_{p})/(x^{\top}t_{n})$. Here, $x,t_{p},t_{n}$ represent embeddings of the output frame, positive (e.g., “Photo”), and negative (e.g., “Game”) prompts, respectively. Higher scores indicate better alignment with real photography.
  • Temporal Consistency: We assess temporal coherence using Motion-S [19] for motion plausibility and WarpSSIM. The latter computes SSIM between a frame and its neighbors warped via optical flow derived from source CG videos.
  • Sim2Real Consistency: We assess semantic preservation using the mean Intersection over Union (mIoU). Specifically, we employ the state-of-the-art perception model [42] (pretrained on Cityscapes) to perform semantic segmentation on the translated frames. We then calculate the mIoU by comparing the predicted segmentation masks against the ground truth semantic labels derived from the source CG engine.
4.3 Quantitative Evaluation.
@@
Table 1: Quantitative evaluations. Top results are colored in Gold, Silver, and Bronze(also applicable to Tabs. 2 and 3.). $\dagger$ denotes modules fine-tuned on the same dataset as DwD. C-T2.5 stands for Cosmos-Transfer 2.5.
@@
C-T2.5 Depth 98.47 92.48 116.50 26.98 38.44
C-T2.5 Edge+Depth 93.97 93.97 112.15 23.71 36.69
C-T2.5 Edge+Seg 98.62 93.97 112.13 23.32 36.62
C-T2.5 Seg 98.85 90.92 119.73 23.81 37.20
Ours 98.94 93.07 119.11 9.30 21.62
4.4 Qualitative Evaluation.